1) The common stock of Zeta Group sells for $42 per share, has a rate of return of 12.2 percent, and a dividend growth rate of 1.8 percent annually. What was the amount of the last annual dividend paid
Answer #1

As per Dividend Discount model  
Price of common stock =Last Dividend*(1+growth)/(Rate-growth)
Last Dividend =Price*(Rate-Growth)/(1+growth) =42*(12.2%-1.8%)/(1+1.8%) =4.29
Answer #2

SOLUTION :


Current price, P0 = 42 ($)

Rate of return, r = 12.2 % = 0.122

Dividend growth rate, g = 1.8% = 0.018


Let $ D0 be the last annual dividend paid.


Now,


D0 (1 + g) = (r - g) * P0

=> D0 * (1 + 0.018) = (0.122 - 0.018) * 42

=> 1.018 D0 = 4.368

=> D0 = 4.368/1.018

=> D0 = 4.29 ($) 


So, amount of the last annual dividend paid = $4.29 (ANSWER).


answered by: Tulsiram Garg